Find the equation of the line passing through the point (3,0,1) and parallel to the planes x+2y = 0 and 3y-z = 0

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
The equation of the line passing through the point (3,0,1) and parallel to the planes x+2y = 0 and 3y-z = 0 is given by
\begin{vmatrix} x-3 &y-0 &z-1 \\ 1&2 &0 \\ 0 & 3 &-1 \end{vmatrix}=0
This gives the equation 2x-y-3z=3